OpenStudy (anonymous):

Using complete sentences, explain how to find the equation of the line, in standard form and slope-intercept form, passing through (3, 6) and (-2, -4). Also compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope-intercept form.

OpenStudy (anonymous):

im jsut gonna tell u ur standard form is simply y=ax+b ur slope form is (y2-y1)/(x2-x1)=a the slope form doesnt give u the b value which is ur y intercept or initial value. :) can u do the problem now? :)

OpenStudy (anonymous):

how do I still do the standard form

OpenStudy (anonymous):

alright well u still have to calculate the slope using that equation which is the slope form :P then take either pt and plug it in with ur a to solve for b and tada :D

OpenStudy (anonymous):

ur a will b 2 :) (6- -4)/(3- -2)=2 y=ax+b 6=2*3+b b=0 :)

OpenStudy (anonymous):

Ok so how do I write the equation for that?

OpenStudy (anonymous):

well y=2x+0= y=2x :)

OpenStudy (anonymous):

How do I do it with slope intercept form